### v3

# The Signals of the Doomsday

Abstract
The measured standard model parameters indicate that we might live in a false Higgs vacuum, though with a very long lifetime. However, small black holes can serve as catalysers and significantly speed up the phase transition. In fact, bubbles of true vacuum might already exist in our universe. We calculate the spectrum of Higgs particles produced by such a bubble, and use event generators to study their decay and subsequent evolution of the decay products to obtain the spectrum of emitted photons and neutrinos as a long-range signature. If the propagation of the bubble walls slows down due to interaction with the surrounding matter and plasma, these signals can reach us before the bubble wall hits us, thus representing the signals of the doomsday.
1 Introduction
It is believed that our universe has so far undergone a number of phase transitions at different energy scales, e.g. GUT, electroweak, QCD… Phase transitions usually introduce drastic changes in the structure of the universe, and if they happen late enough, they would almost certainly be fatal to any existing life forms. Until recently, late-time phase transitions have not been a topic of intensive study (see however [1, 2, 3]).
After the Higgs discovery, it became apparent that a scenario similar to the one outlined in [2] (that the Higgs potential might develop instability with our current universe sitting in the unstable false vacuum) might actually be true [4]. This might not be an immediate problem, since the lifetime of our universe in the false vacuum could be many billions of years. However, as shown in [5, 6, 7, 8, 9, 10, 11], small primordial black holes can play the role of Higgs vacuum decay catalysers and significantly increase the tunneling probability. Depending on the small black hole mass and the exact parameters in the Higgs potential, the tunneling probability could be high enough to produce bubbles of true vacuum within the present lifetime of the universe.
Once a bubble of true vacuum is created, it will expand at a speed close to the speed of light. A bubble wall sweeps through the universe and destroys (or modifies beyond recognition) everything it encounters. If the bubble wall is approaching with the speed of light, then no signal emitted from the wall can reach the observer before the wall, simply because the wall is already moving with the maximal possible speed. However, interaction with surrounding plasma and matter could slow down the wall propagation [12]. For example, if the bubble is created one million light years away from us (somewhere between us and Andromeda) and the wall slows down by only $1$ km/s, then the signal can reach us about three years before the bubble wall. It is then of utmost importance to calculate the characteristics of the possible signals that might come from an approaching bubble.
We investigate here the process in which the Higgs tunnelling can leave a signature in the form of produced particles that can potentially reach us before the bubble. In particular, we calculate the spectrum of Higgs particles created due to the vacuum mismatch inside and outside of the bubble. Then, we use HDecay [13, 14] event generator to calculate the immediate Higgs decay products. Since we are mostly interested in obtaining the long range signature (i.e. emitted photons and neutrinos), we then use Pythia [15] to hadronize the decay products and obtain the final photon and neutrino energy spectra. To quantify the amount of particles that can reach us before the bubble wall, we introduce friction between the wall and the environment, which leads to a terminal velocity slower than the speed of light. We then integrate particle production up to the moment when the terminal velocity is reached and show that it is significant enough for a possible detection.
2 The Higgs vacuum decay
In this section, we review the details of the Higgs vacuum decay in the presence of gravity. The high energy effective Higgs potential has been determined by the two-loop calculations in the standard model (without gravity) as
$V_{\rm SM}(\phi)=\frac{1}{4}\left(\lambda_{*}+b\ln^{2}\frac{\phi}{\phi_{*}}\right)\phi^{4}.$ (2.1)
The standard model range of parameters is $0.01\lesssim\lambda_{*}\lesssim 0$ , $0.1M_{p}\lesssim\phi_{*}\lesssim M_{p}$, where, $M_{p}=\sqrt{\frac{1}{8\pi G}}=2.435\times 10^{18}$GeV, and $b\sim 10^{-4}$ [4, 9, 16]. This potential indicates that we live in an unstable vacuum, though with a very long lifetime [17].
As shown in [10] (for earlier work, see [2]), the inclusion of gravity can significantly modify the dynamics of the electroweak phase transition and drastically shorten the lifetime. We can parametrise the modification by including higher-order operators as
$V(\phi)=V_{\rm SM}+\frac{\lambda_{6}}{6}\frac{\phi^{6}}{M_{p}^{2}}+\ldots$ (2.2)
This potential can have a stable true vacuum and supports a first-order phase transition which happens via the nucleation of bubbles of the new vacuum inside the old vacuum.
Figure 1: This figure shows the Higgs potential $V(\phi)$ in Eq. (2.2). The true vacuum is at $\phi\approx 3.6\times 10^{-2}M_{p}$. The parameters are chosen to be within the standard model, i.e. $b=10^{-4}$, $\lambda=-0.001$, $\phi_{*}=0.5M_{p}$, and $\lambda_{6}=0.34$[16].
We adopt the conventions and follow the calculations in [10]. The nucleation rate in presence of gravity is determined by a bounce solution with Euclidean metric signature (+,+,+,+) with an action
$S_{E}=\int_{\mathcal{M}}\left[-\frac{1}{16\pi G}\mathcal{R}+\left(\frac{1}{2}g^{ab}\partial_{a}\phi\partial_{b}\phi+V(\phi)\right)\right]\sqrt{-g}d^{4}x.$ (2.3)
The spacetime geometry is taken to be spherically symmetric
$ds^{2}=f(r)e^{2\delta(r)}d\tau^{2}+\frac{dr^{2}}{f(r)}+r^{2}(d\theta^{2}+\sin^{2}\theta d\varphi^{2}),$ (2.4)
with
$f(r)=1-\frac{2G\mu(r)}{r},$ (2.5)
where $\mu(r)$ is the mass parameter of a black hole remnant. The phase transition is originally initiated by a black hole of a certain seed mass, $M_{+}$, related to the black hole remnant mass as
$M_{+}=\lim_{r\rightarrow\infty}\mu(r).$ (2.6)
We distinguish here between the black hole seed mass, $M_{+}$, which triggers the phase transition, and the black hole remnant mass, $\mu(r)$, which is a leftover from the seed black hole after some of its energy is invested into the phase transition.
The Higgs field equations of motion in this curved background are
$f\phi^{\prime\prime}+f^{\prime}\phi^{\prime}+\frac{2}{r}f\phi^{\prime}+\delta^{\prime}f\phi^{\prime}-V_{,\phi}=0$ (2.7)
$\mu^{\prime}=4\pi r^{2}\Big{(}\frac{1}{2}f{\phi^{\prime}}^{2}+V\Big{)},\ \ \ \ \delta^{\prime}=4\pi Gr{\phi^{\prime}}^{2},$ (2.8)
where $V_{,\phi}\equiv\partial V/\partial\phi$. The prime and double prime denote spatial derivatives. The black hole horizon is at $r=r_{h}$, which is the solution to $f(r_{h})=0$. We will solve these equations numerically in order to get the function $\phi(r)$. To do this, we start from the horizon with a particular remnant parameter, $\mu_{-}$, at the horizon. The black hole horizon is given as $r_{h}=2G\mu_{-}$. At the horizon, the field $\phi$ satisfies the boundary condition
$\mu(r_{h})=\mu_{-}\text{, }\delta(r_{h})=0$ (2.9)
$\phi^{\prime}(r_{h})=\frac{r_{h}V_{,\phi}(\phi(r_{h}))}{1-8\pi Gr_{h}^{2}V(\phi(r_{h}))}.$ (2.10)
At $r\rightarrow\infty$, the field $\phi$ satisfies $\lim_{r\rightarrow\infty}\phi(r)\rightarrow 0$. We use a shooting method which starts from a special $\phi$ at $r=r_{h}$ and check whether $\phi$ becomes 0 for very large values of $r$. If this condition is satisfied, we have a good solution to the equations of motion.
Figure 2: The Higgs field distribution at the moment when a black hole triggers the false vacuum decay. The inner black hole mass is $1M_{p}$. The radius of the true vacuum bubble at that moment is $2\times 10^{4}M_{p}^{-1}$. The parameters are the same as in Fig. 1.
The Higgs potential $V(\phi)$ from Eq. (2.2) is shown in Fig. 1, while the solution for the Higgs field distribution in the presence of gravity in Fig. 2. Since we have freedom in choosing the exact values of parameters, we follow [16], and for illustration we set $b=10^{-4}$, $\lambda=-0.001$, $\phi_{*}=0.5M_{p}$, and $\lambda_{6}=0.34$.
3 The true Higgs vacuum bubble propagation
As shown in [8, 9, 10], any primordial black hole lighter than $4\times 10^{14}$g at the time of formation would have evaporated by now, and in the absence of new physics beyond the standard model, would have entered the mass range in which it can trigger false vacuum decay (for related work see [18, 19, 20, 21, 22, 23, 24, 25, 26, 27]). When the Higgs field tunnels through the vacuum barrier, a bubble of true vacuum is created and starts expanding. We assume that the geometry around the bubble is not far from a flat spacetime, so that we can neglect gravitational effects. In this case, the pure Higgs field , $\phi(\rho)$, action can be written as
$S(\phi)=\int d^{4}x\Big{(}\frac{1}{2}(\partial_{\mu}\phi)^{2}-V(\phi)\Big{)},$ (3.1)
where we neglected other fields that couple to the Higgs. The equation of motion is
$\partial_{t}^{2}\phi+\nabla^{2}\phi-V^{\prime}(\phi)=0.$ (3.2)
To proceed, we perform a Wick rotation, $t\rightarrow i\tau$, and the equation of motion becomes
$\partial_{\tau}^{2}\phi+\nabla^{2}\phi-V^{\prime}(\phi)=0.$ (3.3)
If we consider an $O(4)$ symmetric solution, this equation can be written as
$\frac{d^{2}\phi_{c}}{d\rho^{2}}+\frac{3}{\rho}\frac{d\phi_{c}}{d\rho}=V^{\prime}(\phi_{c}).$ (3.4)
In this case, $\rho=\sqrt{\tau^{2}+r^{2}}$, while $\phi_{c}(\rho)$ depends only on $\rho$. This equation provides a bounce solution with $\partial_{t}\phi(t=0)=0$ and $\phi(t=0,\vec{x})=\phi_{c}(\tau=0,\vec{x})$ as
$\phi(t,\vec{x})=\phi_{c}(\rho=\sqrt{r^{2}-t^{2}}).$ (3.5)
If we plot $r^{2}-t^{2}=\rho^{2}=$constant, we will find that the minimal $r$ happens at $t=0$. This is the moment when the bubble bounces back. Thus, the Higgs bubble is collapsing for $t<0$, bounces at $t=0$, and is expanding for $t>0$. The region $r>t$ is still in the false vacuum, while the $r<t$ region is in the true vacuum. So in this parametrization, nucleation happens at $t=0$, and then the true vacuum bubble expands. We simplify the problem by applying the thin wall approximation
$\phi_{c}(\rho)=\left\{\begin{array}[]{lr}v&\text{, for $\rho>R$}\\ v_{1}&\text{, for $\rho<R$}\end{array}\right$., (3.6)
where $v$ and $v_{1}$ are the expectation values of the Higgs field in the false and true vacuum respectively, while $R$ is the radius of the bubble.
Creation of particles during the first-order phase transitions has been studied previously (e.g., [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]). In principle, particles can be produced via many different mechanisms. For our purpose, we will use the vacuum mismatch method [33, 39] which can be easily formulated in Minkowski (as opposed to Euclidean) space.
4 Particle production due to vacuum mismatch
The mismatch of vacua is usually accompanied by particle production. In our case, the false Higgs vacuum is outside the bubble, while the true vacuum is inside. As the bubble expands through space, the Higgs field vacuum abruptly changes from false to true. Bogolyiubov transformations then indicate that this vacuum mismatch will lead to the Higgs particle production. Since the Higgs particle is very heavy in the new vacuum, its decay products will be very energetic. For simplicity, we assume that the Higgs field tunnels to the true vacuum state homogeneously. This should be a good enough approximation since particles are produced locally, i.e. the relevant phase transition length scale is smaller than the size of the bubble at virtually any time after the initial bubble nucleation. Here, for the relevant length scale we take the width of the potential barrier in Fig. 1 ($\sim[10^{-2}M_{p}]^{-1}$), which is also the scale of the Higgs expectation value in the true vacuum.
The fluctuations of the Higgs field, $\phi$, around the background, $\phi_{c}$, can be written as $\phi=\phi_{c}+h$, where $h$ is the fluctuation which satisfies the equation of motion
$\partial_{\tilde{\tau}}^{2}h+\nabla^{2}h-V^{\prime\prime}(\phi_{c})h=0.$ (4.1)
We again considered only the Higgs part of the Lagrangian and neglected other fields that couple to the Higgs. In this case the equation of motion (4.1) can be simplified to
[…]
$\partial_{\tilde{\tau}}^{2}h+\nabla^{2}h-\mu^{2}h=0\text{, for $\tilde{\tau}>\tilde{\tau}^{*}$}$, (4.3)
where $\tilde{\tau}^{*}$ is the characteristic time scale for the duration of the transition between the vacua. Since the number density of the produced particles crucially depends on this constant (see Eqs. (4.7) and (4.9)) we have to choose it in a meaningful way. We will set $\tilde{\tau}^{*}=-R_{0}$, where $R_{0}$ is the radius of the bubble at the time of nucleation. The justification for this choice is that this is the magnitude of the proper acceleration of the shell, i.e. $a=1/R_{0}$. The process of particle creation due to the vacuum mismatch is in this context closely related to Unruh radiation, which is determined by the magnitude of the proper acceleration. We note here that this magnitude is a Lorentz invariant quantity, $a=\sqrt{a_{\mu}a^{\mu}}$, which can be calculated from the equation of the bubble motion, $r^{2}-t^{2}=R_{0}^{2}$ (see Appendix-A [10]). As such, it is not equal to the coordinate acceleration that an outside observer would see. The proper acceleration is rather parametrized in terms of the proper time for an observer who is momentarily at rest with respect to the bubble wall. Since in our case $a=1/R_{0}$, wall expansion is a motion with a constant proper acceleration, and particles are produced continuously. Alternatively, we could take $\tilde{\tau}^{*}=-\eta^{-1}$, where $\eta\sim 10^{-2}M_{p}$ is the characteristic energy scale for the phase transition (i.e. the Higgs vacuum expectation value in the new vacuum, and also the width of the potential barrier). This choice will yield significantly more produced particles, however it lacks connection with the Unruh effect. We therefore go with the more conservative choice $\tilde{\tau}=-R_{0}$.
The solution for $h$ in Eq. (4.2) can be written as a combination of mode functions $g_{k}$, which satisfy $\triangledown^{2}g_{k}=-k^{2}$
[…]
Here, $\omega_{+}=\sqrt{\mu^{2}+k^{2}}$ and $\omega_{-}=\sqrt{M^{2}+k^{2}}$, while $M=\sqrt{V^{\prime\prime}(v)}$ and $\mu=\sqrt{V^{\prime\prime}(v_{1})}$ are the masses of the Higgs field in the false and true vacuum regions, respectively. Since $g_{k}$ and $\partial_{\tilde{\tau}}g_{k}$ must be continuous at $\tilde{\tau}=\tilde{\tau}^{*}=-R_{0}$, for $A_{k}$ and $B_{k}$ we get
$A_{k}$ $=$ $\frac{1}{2\omega_{+}}(\omega_{+}+\omega_{-})e^{-(\omega_{+}-\omega_{-})\tilde{\tau}^{*}}$ (4.7)
$B_{k}$ $=$ $\frac{1}{2\omega_{+}}(\omega_{+}-\omega_{-})e^{(\omega_{+}+\omega_{-})\tilde{\tau}^{*}}$. (4.8)
The particle creation spectrum is obtained from the Bogoliubov transform [39]
$N_{k}=\frac{B_{k}^{2}}{A_{k}^{2}-B_{k}^{2}}=\left[\frac{(\omega_{+}+\omega_{-})^{2}}{(\omega_{+}-\omega_{-})^{2}}e^{4\omega_{+}R_{0}}-1,\right]^{-1}$ (4.9)
Figure 3: Number density of Higgs particles as a function of their momenta created due to the vacuum mismatch in the Higgs vacuum decay. The units are given in terms of the Planck mass $M_{p}$.
Fig. 3 shows the number density of the produced Higgs particles per momentum mode per unit volume, $N_{k}=dN/(dVd^{3}\vec{k})$, as a function of their momenta, $k$, given in Planck units. The value of the Higgs’ mass in the false vacuum (where we currently live) is $M=125.09\,\text{GeV}$ [40]. For the choice of parameters we used in Fig. 1, the Higgs mass in the true vacuum is $\mu=7.16\times 10^{-4}M_{p}$. We emphasize that we do not have much freedom here, and this value might change by only one order of magnitude if we want to keep the values for $b,\lambda$, and $\phi_{*}$ within the standard model. For the same choice of parameters the bubble radius at the time of nucleation is $R_{0}=2\times 10^{4}M_{p}^{-1}$.
We can estimate the energy density of Higgs particles inside the bubble (in the new vacuum) at the time of creation as
$\mathcal{E}$ $=$ $\int_{0}^{\infty}\omega_{+}N_{k}\frac{d^{3}k}{(2\pi)^{3}}$ (4.10)
$=$ $\int_{0}^{\infty}\frac{\sqrt{\mu^{2}+k^{2}}}{\frac{(\sqrt{\mu^{2}+k^{2}}+\sqrt{M^{2}+k^{2}})^{2}}{(\sqrt{\mu^{2}+k^{2}}-\sqrt{M^{2}+k^{2}})^{2}}e^{4\sqrt{\mu^{2}+k^{2}}R_{0}}-1}$
$=$ $2.4\times 10^{-42}M_{p}^{4}=8.6\times 10^{31}\,\text{GeV}^{4}.$
For any bubble of macroscopic volume, this is a very significant total energy which is ultimately transferred into the decay products. Thus, the bubble of the true vacuum should be a very shiny object.
5 Higgs decay into photons and neutrinos
A complete treatment of the Higgs decay is very complicated under these circumstances. Inside the bubble the Higgs field is in a new vacuum, so it is heavy, while outside is in the old vacuum, and it is light. A heavy Higgs cannot exist outside of the bubble, and it has to decay practically instantaneously. The decay products have to propagate outside in the old vacuum to reach a distant observer. Strictly speaking, one would have to model how the heavy Higgs propagates through the barrier and decays along the way. To simplify the task, we just assume that the Heavy Higgs decays instantaneously at the bubble wall. Ultimately, we want the spectrum of photons and neutrinos as a long-range observational signature. For that purpose, we will use the event generators that simulate the Higgs decay and subsequent evolution of the decay products. Generators like HDecay are convenient for our purpose since they evolve coupling constants with energy and can simulate heavy Higgs decay.
To simplify the analysis, we set $k=0$ in the definitions of $\omega_{+}$ and $\omega_{-}$, i.e. we consider the decay of Higgs particles at rest. Using this assumption, we calculate the number of Higgs particles produced due to vacuum mismatch from equation (4.9) for the specific value of the Higgs’s mass in true vacuum of $\mu=7.16\times 10^{-4}M_{p}$, resulting in $N_{k}(k=0)\simeq 1.365\times 10^{-25}$ (per Planck volume). Furthermore, we assume that the standard model (SM) remains valid up to the true vacuum range at $\phi\approx 3.6\times 10^{-2}M_{p}$. Our analysis includes only the leading-order contributions to Higgs decay to the fermions, disregarding higher-order corrections. This approximation is justified within the framework of the effective field theory, where the higher-order terms are highly suppressed at the mass scales under consideration.
To obtain the branching ratios and the total decay width of the SM Higgs particle at a true vacuum, we use HDECAY. We use only the tree level diagrams and switch off the electroweak corrections for the simulations because they become unstable beyond the 1 TeV range. For the rest of the input parameters, we follow the instructions given in [13] without any changes.
The output results for the branching ratios into fermions and bosons, as well as the total decay width, are listed in Tables 1 and 2, respectively.
Table 1: Branching ratios for the Higgs decays into fermions.
$M_{H}$ (GeV) $BR({b\bar{b}})$ $BR({c\bar{c}})$ $BR({s\bar{s}})$ $BR({t\bar{t}})$ $BR({\tau^{+}\tau^{-}})$ $BR({\mu^{+}\mu^{-}})$
$1.75\times 10^{15}$ $6.08\times 10^{-31}$ $2.85\times 10^{-32}$ $1.72\times 10^{-34}$ $1.57\times 10^{-27}$ $1.35\times 10^{-30}$ $4.77\times 10^{-33}$
Table 2: Branching ratios for the Higgs decays into bosons and the total decay width.
$M_{H}$ (GeV) $BR(GG)$ $BR(\gamma\gamma)$ $BR(Z\gamma)$ $BR(WW)$ $BR(ZZ)$ Width (GeV)
$1.75\times 10^{15}$ $1.08\times 10^{-04}$ $46\times 10^{-07} 9.9$$90\times 10^{-07}$ $6.95\times 10^{-01}$ $3.49\times 10^{-01}$ $2.30\times 10^{39}$
Note that the Higgs decay width can become larger than the Higgs mass, which indicates that the Higgs behaves more like a resonance than a well-defined particle at these scales, as argued in [41]. The output data consisting of the Higgs mass, branching ratios and the Higgs decay width, generated from HDecay, were used as input parameters into Pythia for hadronization (formation of mesons and baryons) and final-state decays. In Pythia, we successfully simulated hadronization and decay for various channels up to a Higgs mass of $1\times 10^{7}$ GeV without encountering simulation errors. We then extrapolated these findings (number of photons and neutrinos) using the logarithmic extrapolation method, to the true vacuum region at a Higgs mass of $7.16\times 10^{-4}M_{p}=1.75\times 10^{15}$ GeV, which provided the most accurate predictions when compared with theoretical decay width calculations for each decay channels. After extrapolation, we finally get the observational signature in terms of the final-state particles.
In the final state, the most abundant particles are photons and neutrinos, while stable particles such as electrons, muons, protons, neutrons, and other heavy baryons are produced in significantly smaller quantities. These final-state particles primarily originate from Higgs decays into quarks, which subsequently hadronize. Since photons and neutrinos vastly outnumber the stable particles, they represent the primary observational signature of Higgs vacuum decay, allowing us to neglect the contribution from stable particles. They anyway represent the desired long-range signature of the bubble formation.
Figures 4 and 5 show the spectra of photons and neutrinos respectively, i.e. their number densities as a function of their energies, originated in decay of $N_{k}(k=0)\simeq 1.365\times 10^{-25}$ Higgs particles per unit volume produced by a bubble of true vacuum.
Figure 4: The spectrum of photons created by a bubble of true vacuum due to the Higgs decay.
Figure 5: The spectrum of neutrinos created by a bubble of true vacuum due to the Higgs decay.
During the expansion of the bubble, vacuum is continuously changing from an old one to a new one, so particles are produced continuously in this process. Since the plots in 4 and 5 are actually particle number densities given in Planck units, to get the total number of produced particles (as a function of energy) we have to integrate over the volume of the bubble.
We also note that the spectra in Figures 4 and 5 show the particle number density at the site of creation. To obtain the observed spectra on Earth, we have to propagate these particles through expanding space. The corresponding particle flux (number of particles per unit area per unit time) would therefore be diluted by a factor of $1/[4\pi d^{2}(1+z)]$, where $d$ is the physical distance between the source and the observer, while $z$ is the redshift of the source. The factor $(1+z)$ appears due to the relativistic time delay because particles hit the sphere with the radius $d$ less frequently (two particles emitted $\Delta t$ apart will be measured $(1+z)\Delta t$ apart. Accordingly, we have to correct the particle energies by a factor of $1+z$ because the individual particle’s energies are redshifted by that factor.
6 Relativistic bubble-wall dynamics in a viscous medium and terminal velocity
If the bubble wall moves with the speed of light, then no signal can overtake it to warn us about the impending doom. However, a bubble interacts with its environment and can reach a terminal velocity that is less than the speed of light. Previously, this topic has been studied extensively in the literature, with works addressing bubble wall velocities, hydrodynamical constraints, and friction mechanisms in various contexts[42, 43, 44, 45, 46, 47, 48]. We begin by noting that a bubble wall propagating through matter or radiation (where some of it could be produced by the bubble itself) experiences friction due to particle interactions at its surface. Understanding this damping is crucial, as friction reduces the wall’s acceleration and sets its terminal velocity, thereby shaping the overall dynamics and determining the efficiency of associated particle production.
6.1 Derivation of the equation of motion
In this section, we present a step-by-step derivation of the relativistic equation of motion for a spherical thin bubble wall, incorporating the driving vacuum pressure, the Laplace (curvature) pressure, and a linear frictional damping that arises from interactions with the surrounding particles.
The energy per unit area of the static wall is its surface tension,
$E_{\rm rest}=\sigma,$ (6.1)
which under a Lorentz boost becomes
$E_{\rm wall}=\sigma\,\gamma(v),\qquad\gamma(v)=\frac{1}{\sqrt{1-v^{2}}},$ (6.2)
while the corresponding momentum per unit area is
$\Pi_{\rm wall}=E_{\rm wall}\,v=\sigma\,\gamma(v)\,v.$ (6.3)
Here, $v$ is the wall velocity. Taking the time derivative of equation (6.3) yields the inertial response of the wall:
$\frac{d}{dt}\Pi_{\rm wall}=\sigma\,\frac{d}{dt}[\gamma(v)\,v]=\sigma\,\gamma^{3}(v)\,\frac{dv}{dt},$ (6.4)
where we have used
$\frac{d}{dt}[\gamma v]=\bigl{[}\gamma+v\,\gamma^{\prime}(v)\bigr{]}\,\frac{dv}{dt}=\gamma^{3}(v)\,\frac{dv}{dt}.$ (6.5)
This result encapsulates the relativistic increase in inertia as $v\to 1$.
For a spherical wall of radius $R(t)$, the net outward force per unit area combines the vacuum-pressure jump, the Laplace pressure, and the frictional pressure:
$F_{\rm net}\;=\;\underbrace{\Delta V}_{\text{vacuum drive}}\;-\;\underbrace{\frac{2\sigma}{R(t)}}_{\text{Laplace curvature}}\;-\;\underbrace{\eta\,\gamma(v)\,v}_{\text{linear drag}}\,.$ (6.6)
where $\Delta V=V_{\text{false}}-V_{\text{true}}$ is the pressure difference (latent heat), while $\eta$ is the friction coefficient. Balancing momentum change against the net force gives the spherical thin-wall equation of motion (EOM):
$\sigma\,\gamma^{3}(v)\,\frac{dv}{dt}=\Delta V\;-\;\frac{2\sigma}{R(t)}\;-\;\eta\,\gamma(v)\,v.$ (6.7)
Equation (6.7) is the relativistic friction-damped spherical wall equation. It shows that as the wall accelerates, both the relativistic inertia (via $\gamma^{3}$) and the frictional drag (via $\gamma v$) increase, while curvature adds an extra restoring pressure $\sim 2\sigma/R$ that is important at early times (small $R$) and fades at late times ($R\to\infty$).
At the instant of terminal balance, the acceleration vanishes ($dv/dt=0$). Equation (6.7) then gives the curvature-modified balance condition
$0=\Delta V\;-\;\frac{2\sigma}{R}\;-\;\eta\,\gamma(v)\,v\quad\Longleftrightarrow\quad\Delta V_{\rm eff}(R)\;\equiv\;\Delta V-\frac{2\sigma}{R}=\eta\,\gamma(v)\,v.$ (6.8)
In the late-time, large-radius limit $R\to\infty$ one recovers the planar terminal relation $\Delta V=\eta\,\gamma v$.
6.2 Terminal Velocity
Calculating the terminal velocity of the wall explicitly requires the knowledge of the frictional coefficient, $\eta$, in a given medium, which is possible in some simple situations, but not in the general case.
The bubble wall reaches terminal velocity when the driving force balances the friction force. The equation of motion for the wall (neglecting its curvature) is:
$\Delta V=\eta\gamma v.$
For highly relativistic walls ($\gamma\gg 1$, $v\approx 1$), this simplifies to:
$\Delta V\approx\eta\gamma.$
Thus, the terminal Lorentz factor is:
$\gamma_{\text{term}}\approx\frac{\Delta V}{\eta}.$
The terminal velocity is then:
$v_{\text{term}}\approx 1-\frac{1}{2\gamma_{\text{term}}^{2}}.$
The friction coefficient $\eta$ can be calculated from microphysics. For a scalar field $\phi$ coupled to other particles (e.g., fermions) with mass $m\sim g\phi$, the friction can be estimated as:
$\eta\sim g^{2}\Gamma,$
where $g$ is the coupling constant, while $\Gamma$ is the interaction rate of particles with the wall. Suppose that the wall is moving in a hot dense plasma. In thermal equilibrium, $\Gamma\sim T^{4}$ (temperature of the plasma)[49], so
$\eta\sim g^{2}T^{4}.$
For the order of magnitude estimate, within the standard model, we can set $g\approx 0.1$. We can also generically set $\Delta V\sim T^{4}$. This gives
$v_{\text{term}}\approx 1-\frac{1}{2\gamma_{\text{term}}^{2}}=0.995c.$
This estimate may be acceptable in the early universe or in cases where the bubble is produced inside stars. However, in more complicated situations where the bubble is created in the late universe in vacuum, and encounters stars, planets, and diluted interstellar and intergalactic gas during its expansion, more detailed numerical simulations are warranted. After all, the bubble produces a large number of very high energy particles, so it is literally engulfed in its own plasma. Thus, precise calculations of the terminal velocity will be studied elsewhere, and we will instead consider several conservative values for illustration.
6.3 Proper acceleration of a spherical bubble wall in a viscous medium
Since particle production we are describing here is a variant of the Unruh effect [50, 51] driven by the proper acceleration of the wall, we will derive now the invariant (proper) acceleration of a relativistic spherical thin wall when both curvature and linear friction are present. The result provides the quantity that directly controls vacuum-mismatch particle production.
We start from the bubble wall equation of motion given in Eq. (6.7)
$\sigma\,\gamma^{3}(v)\,\frac{dv}{dt}=\Delta V-\frac{2\sigma}{R(t)}-\eta\,\gamma(v)\,v.$ (6.9)
We introduce the wall’s proper time $\tau$ and the rapidity $y(\tau)$ via
$v=\tanh y,\qquad\gamma=\cosh y,\qquad\gamma v=\sinh y,\qquad\frac{dt}{d\tau}=\gamma,\qquad\frac{dR}{d\tau}=\gamma v=\sinh y.$ (6.10)
The invariant (proper) acceleration is defined by
$\alpha(\tau)\;\equiv\;\frac{dv}{d\tau}\,\gamma^{2}\;=\;\gamma^{3}\,\frac{dv}{dt}\;=\;\frac{dy}{d\tau}.$ (6.11)
Dividing (6.9) by $\sigma$ and using (6.11) yields the proper-time form:
$\frac{dy}{d\tau}$ $=\underbrace{\frac{\Delta V}{\sigma}}_{\displaystyle A}\;-\;\underbrace{\frac{2}{R(\tau)}}_{\text{curvature}}\;-\;\underbrace{\frac{\eta}{\sigma}}_{\displaystyle B}\,\sinh y,$ (6.12)
$\frac{dt}{d\tau}$ $=\cosh y,\qquad\frac{dR}{d\tau}=\sinh y.$ (6.13)
Equivalently, the proper acceleration is explicitly
$\alpha(\tau)=A\;-\;\frac{2}{R(\tau)}\;-\;B\,\sinh y(\tau)\,,\qquad A\equiv\frac{\Delta V}{\sigma},\quad B\equiv\frac{\eta}{\sigma}\;$}. (6.14)
We take the nucleation initial data
$R(0)=R_{0},\qquad y(0)=0\ (v=0),\qquad t(0)=0.$ (6.15)
From (6.14) the initial proper acceleration is
$\alpha(0)=A-\frac{2}{R_{0}}.$ (6.16)
In the thin-wall nucleation one has the standard relation $A\simeq 3/R_{0}$, yielding
$\alpha(0)\;\simeq\;\frac{1}{R_{0}},$ (6.17)
which is just the acceleration $a$ we used in Section 4.
For $\tau\ll R_{0}$ the motion is mildly relativistic. Using $y(\tau)=\alpha(0)\tau+\mathcal{O}(\tau^{2})$, we have for small $y$ $\sinh y\simeq y,$ hence
$\sinh y\simeq\alpha(0)\,\tau,\qquad R(\tau)\simeq R_{0}+\int_{0}^{\tau}\sinh y\,d\tau^{\prime}\;\simeq\;R_{0}+\frac{\alpha(0)}{2}\,\tau^{2}.$ (6.18)
Then by inserting (6.18) into (6.14) we get the friction term:
$B\,\sinh y(\tau)\simeq-B\,\alpha(0)\,\tau(\text{linear in}\,\tau).$ (6.19)
To simplify the curvature term we expand $1/R(\tau)$ about $R_{0}$:
$\frac{1}{R(\tau)}=\frac{1}{R_{0}+\frac{\alpha(0)}{2}\tau^{2}}=\frac{1}{R_{0}}\,\frac{1}{1+\frac{\alpha(0)}{2R_{0}}\tau^{2}}\simeq\frac{1}{R_{0}}\left(1-\frac{\alpha(0)}{2R_{0}}\tau^{2}\right),$ (6.20)
so
$\frac{2}{R(\tau)}\simeq-\frac{2}{R_{0}}+\frac{\alpha(0)}{R_{0}^{2}}\,\tau^{2}.$ (6.21)
$\alpha(\tau)\simeq\underbrace{\Big{(}A-\frac{2}{R_{0}}\Big{)}}_{\alpha(0)}\;-\;B\,\alpha(0)\,\tau\;+\;\frac{\alpha(0)}{R_{0}^{2}}\,\tau^{2}\;+\;\mathcal{O}(\tau^{3}).$ (6.22)
With the thin–wall nucleation relation $A\simeq 3/R_{0}$, the initial proper acceleration is
$\alpha(0)=A-\frac{2}{R_{0}}\;\simeq\;\frac{1}{R_{0}}\equiv a.$ (6.23)
Inserting $\alpha(0)=1/R_{0}$ into the early–time series $\alpha(\tau)\simeq\alpha(0)-B\,\alpha(0)\tau+\alpha(0)\tau^{2}/R_{0}^{2}+\mathcal{O}(\tau^{3})$ yields
$\alpha(\tau)\;\simeq\;\frac{1}{R_{0}}\;-\;\frac{B}{R_{0}}\,\tau\;+\;\frac{\tau^{2}}{R_{0}^{3}}\;+\;\mathcal{O}\!\big{(}\tau^{3}\big{)},$ (6.24)
valid in the mildly–relativistic regime $\tau\ll R_{0}$ (so $y\ll 1$) and for weak friction in the sense $BR_{0}\ll 1$. In this expansion, the linear term $(B/R_{0})\tau$ comes from the viscous drag $B\,\sinh y$, while the quadratic term $\tau^{2}/R_{0}^{3}$ originates from the curvature relaxation in $2/R(\tau)$ via $R(\tau)\simeq R_{0}+\tfrac{1}{2}\alpha(0)\tau^{2}$. Higher–order pieces $\mathcal{O}(\tau^{3})$ include small corrections from the next orders in the $\sinh y$ and $1/R$ expansions and are negligible under the same conditions.
Treating (6.24) as a quadratic polynomial in $\tau$, $\alpha(\tau)\approx a\,\tau^{2}+b\,\tau+c$ with $a=1/R_{0}^{3}$, $b=-(B/R_{0})$ and $c=1/R_{0}$, this parabola (with $a>0$) minimum occurs at
$\tau_{\star}=\frac{-b}{2a}=\frac{(B/R_{0})}{2/R_{0}^{3}}={\;\frac{1}{2}\,B\,R_{0}^{2}\;},\qquad\alpha_{\min}=c-\frac{b^{2}}{4a}={\;\frac{1}{R_{0}}-\frac{1}{4}\,B^{2}\,R_{0}\;}.$ (6.25)
$\tau_{\star}$ is the early-time proper time (measured on the wall) at which the quadratic Taylor approximation for the proper acceleration reaches its local minimum. Physically, $\tau_{\star}$ marks the instant where curvature’s quadratic “recovery” just balances the linear drag trend, and $\alpha_{\min}$ is the corresponding (minute) reduction of the net drive. The impact on instantaneous vacuum–mismatch production is therefore negligible at early times. Beyond this local regime, as $y$ grows the friction term $B\sinh y$ dominates and the full evolution drives $\alpha(\tau)\to 0^{+}$ toward terminal balance; thus $\alpha_{\min}$ here is an early–time feature of the quadratic approximation, not a global minimum of the complete dynamics.
We differentiate (6.14) and use (6.13):
$\frac{d\alpha}{d\tau}$ $=\frac{d}{d\tau}\!\left(A-\frac{2}{R}-B\sinh y\right)=\frac{2}{R^{2}}\,\frac{dR}{d\tau}\;-\;B\,\cosh y\,\frac{dy}{d\tau}$
$=\frac{2\,\sinh y}{R^{2}}\;-\;B\,\cosh y\,\alpha.$ (6.26)
In the planar, small–rapidity limit ($R\to\infty$, $\cosh y\simeq 1$) one has $d\alpha/d\tau\simeq-B\,\alpha$, so $\alpha(\tau)\simeq\alpha(0)e^{-B\tau}$. For spherical walls, the positive curvature term slows this decay; we therefore use $\tau_{\rm fric}\equiv 1/B=\sigma/\eta$ as a characteristic timescale for approaching terminal balance. From this we get
$\tau_{term}=1/B=\frac{\sigma}{\eta},$ (6.27)
which defines the characteristic time scale for the terminal velocity to be reached. As $R\to\infty$ the curvature term disappears and (6.14) reduces to
$\alpha(\tau)\;\to\;A-B\,\sinh y(\tau).$ (6.28)
The terminal state is reached when $\alpha\to 0$, i.e.
$\sinh y_{\rm term}=\frac{A}{B}\quad\Longleftrightarrow\quad\gamma_{\rm term}\,v_{\rm term}=\frac{\Delta V}{\eta}.$ (6.29)
The quantity $\tau_{\rm term}$ from (6.27) is important because particles will be produced predominantly up to this moment, when both the proper and coordinate accelerations become zero. In fact, particles might be produced after that moment via some other mechanism, but not due to this vacuum mismatch related to the Unruh effect, which requires a non-zero proper acceleration.
7 Modified Higgs particles production due to the friction
With the friction dynamics in hand, we now estimate how many Higgs quanta are produced in the regime before the terminal velocity is reached. Throughout we work in Planck units, $G=\hbar=c=1$. The relevant parameters taken from our choice of the potential in the previous sections are
$\Delta V=5.869\times 10^{-12}\;M_{p}^{4},\quad R_{0}=2.0\times 10^{4}\;M_{p}^{-1},\quad\mu=7.16\times 10^{-4}\;M_{p},$
$\sigma=5.8680\times 10^{-8}\;M_{p}^{3},\quad M=125.09\,\text{GeV}.$ (7.1)
We use again dimensionless drive and friction ratios
$A\equiv\frac{\Delta V}{\sigma},\qquad B\equiv\frac{\eta}{\sigma},\qquad A=1.000170415815\times 10^{-4}\;M_{p},$ (7.2)
and parametrize the wall by its proper time $\tau$ and rapidity $y(\tau)$:
$v=\tanh y,\quad\gamma=\cosh y,\quad\gamma v=\sinh y,\qquad\frac{dt}{d\tau}=\gamma,\quad\frac{dR}{d\tau}=\sinh y.$ (7.3)
The proper acceleration that controls particle production is
$\alpha(\tau)\;\equiv\;\frac{dy}{d\tau}\;=\;A\;-\;\frac{2}{R(\tau)}\;-\;B\,\sinh y(\tau),$ (7.4)
with nucleation data
$R(0)=R_{0},\qquad y(0)=0\ (v=0),\qquad t(0)=0.$ (7.5)
Note again that $\alpha(0)=A-2/R_{0}\simeq 1/R_{0}$ under the thin-wall relation $A\simeq 3/R_{0}$, and that curvature $2/R$ initially suppresses the drive but becomes negligible as $R$ grows.
Using the proper acceleration $\alpha(\tau)$ from (7.4) we explicitly integrate particle production up to $\tau=\tau_{\rm term}$. Since both the terminal velocity, $v_{\rm term}$, and the time scale $\tau_{\rm term}$ depend on the friction coefficient $\eta$, for illustration, we will consider a few values of $v_{\rm term}$ for which we will calculate $\tau_{\rm term}$ and perform numerical integration.
At proper time $\tau$, we generalize the instantaneous zero-momentum mode occupancy calculated in eq. (4.9) using the vacuum mismatch procedure
$N_{k=0}(\tau)=\left[\frac{(\omega_{+}+\omega_{-})^{2}}{(\omega_{+}-\omega_{-})^{2}}e^{\frac{4\omega_{+}}{\alpha(\tau)}}-1,\right]^{-1},$ (7.6)
where we replaced $R_{0}=1/\alpha(0)$ with $R_{0}=1/\alpha(\tau)$ to take into account that the proper acceleration decreases due to friction. Also, after setting $k=0$, we have $\omega_{+}=\mu$ and $\omega_{-}=M$. This $N_{k=0}(\tau)$ vanishes automatically when $\alpha(\tau)=0$ (terminal balance). During a small proper-time step $d\tau$, the shell of volume swept by the wall is $4\pi R^{2}\,v\,dt=4\pi R^{2}\,\sinh y\,d\tau$. Multiplying this new volume by the instantaneous number per unit volume $N_{k=0}(\tau)$ gives the increment in the number of particles produced. Therefore, the instantaneous number production is the occupancy times the area and the lab-time increment:
$dN=N_{k=0}(\tau)\;4\pi R(\tau)^{2}\;v(\tau)\,dt=N_{k=0}(\tau)\;4\pi R(\tau)^{2}\;\sinh y(\tau)\,d\tau,$ (7.7)
so that the accumulated yield satisfies
$\frac{dN_{\rm tot}}{d\tau}=N_{k=0}(\tau)\;4\pi R(\tau)^{2}\;\sinh y(\tau),\qquad N_{\rm tot}(0)=0.$ (7.8)
$N_{\rm tot}(\tau)$ is the cumulative number produced up to time $\tau$ by integrating those increments beginning at nucleation. At the instant of nucleation $(\tau=0)$ no time has elapsed and no volume has been swept, so by definition the accumulated total is zero: $N_{\rm tot}(0)=0.$ $N_{k=0}(0)$ is, on the other hand, an instantaneous occupation per unit volume; we still need a finite swept volume (over a finite $d\tau$) to accumulate a non-zero total.
Numerical procedure
We discretize proper time $\tau$ and integrate the coupled system $\{y(\tau),t(\tau),R(\tau),N_{\rm tot}(\tau)\}$ defined by (7.3)–(7.8) using a classical fourth-order Runge–Kutta (RK4) method with adaptive steps. The state is initialized at nucleation by $y(0)=0$ $(v=0)$, $t(0)=0$, $R(0)=R_{0}$, and $N_{\rm tot}(0)=0$. At each substep we evaluate the raw proper acceleration
$\alpha_{\rm raw}(\tau)=A-\frac{2}{R(\tau)}-B\,\sinh y(\tau).$
If $\alpha_{\rm raw}\leq 0$ (force balance), we impose the terminal-balance kinematic clamp $dy/d\tau=0$ so the rapidity ceases to grow; in the same branch we set $N_{k=0}(\tau)=0$ in (7.6), so production halts consistently with the $\alpha\to 0^{+}$. Otherwise we update $y,t,R$ via
$\frac{dy}{d\tau}=\alpha_{\rm raw},\qquad\frac{dt}{d\tau}=\cosh y,\qquad\frac{dR}{d\tau}=\sinh y,$
and accumulate the produced quanta using
$\frac{dN_{\rm tot}}{d\tau}=N_{k=0}(\tau)\,4\pi\,R(\tau)^{2}\,\sinh y(\tau),\qquad N_{k=0}(\tau)=\left[\frac{(\omega_{+}+\omega_{-})^{2}}{(\omega_{+}-\omega_{-})^{2}}e^{\frac{4\omega_{+}}{\alpha_{raw}}}-1,\right]^{-1}.$
Stability and stiffness handling: Because $e^{4\mu/\alpha_{\rm raw}}$ can be enormous when $\alpha_{\rm raw}$ is small, we evaluate (7.6) with a numerically stable asymptotic branch,
$N_{k=0}(\tau)\;\simeq\;\left[\frac{(\omega_{+}+\omega_{-})^{2}}{(\omega_{+}-\omega_{-})^{2}}\right]^{-1}e^{-4\mu/\alpha_{\rm raw}}\quad\text{when}\quad\frac{4\mu}{\alpha_{\rm raw}}\gtrsim 200,$
which prevents overflow while preserving accuracy. The proper-time step is adapted according to the local drive magnitude,
$d\tau\;\sim\;\frac{0.05}{\max\!\big{(}|\alpha_{\rm raw}|,10^{-6}\big{)}},$
then clamped to a conservative band (e.g. $20\leq d\tau\leq 800$ in Planck units) to keep $\Delta y$ modest and ensure smooth $R$ growth. This adaptivity is crucial in two stiff regimes: (i) very early times, when curvature $2/R$ is large, and (ii) near terminal, where $\alpha_{\rm raw}\to 0^{+}$ and the evolution slows.
Stopping criterion and diagnostics. The integration stops when the coordinate time first satisfies $t(\tau)=\tau_{\rm term}$ for the scenario. For accuracy control, we repeat each run with the initial $d\tau$ halved and demand relative changes $\lesssim 10^{-4}$ in $\{\tau_{\rm term},R_{\rm fin},N_{\rm tot}^{\rm(int)}\}$. As additional consistency checks we verify: (a) the differential inequality $d\alpha/d\tau\leq-B\,\alpha$ (hence $\alpha(\tau)\leq\alpha(0)e^{-B\tau}$) is respected pointwise; (b) monotonicity of $R(\tau)$; and (c) that $N_{k=0}(\tau)$ vanishes whenever $\alpha_{\rm raw}\leq 0$. To illustrate, we consider three ultra-relativistic terminal deficits $\delta\equiv 1-v_{\rm term}$ and use the given friction pairs $(\eta,\tau_{\rm term})$ for each scenario; this fixes $B=\eta/\sigma$ and thus the proper-time dynamics via (7.4)–(7.8). At $t=\tau_{\rm term}$ we report: the final proper time $\tau_{\rm fin}$, radius $R_{\rm fin}$, and the integrated yield $N_{\rm tot}^{\rm(int)}$.
The friction parameter $\eta$ is fixed self-consistently by the terminal balance condition
$\gamma_{\rm term}\,v_{\rm term}\;=\;\frac{\Delta V}{\eta},$ (7.9)
which follows from setting the driving pressure $\Delta V$ equal to the drag force. For a chosen ultra-relativistic deficit $\delta=1-v_{\rm term}$ one computes $v_{\rm term}=1-\delta$ and $\gamma_{\rm term}=1/\sqrt{1-v_{\rm term}^{2}}$, so that
$\eta(\delta)\;=\;\frac{\Delta V}{\gamma_{\rm term}\,v_{\rm term}}.$ (7.10)
The integration limit $\tau_{\rm term}$ is given in (6.27). This procedure determines the appropriate $\eta$ value for each scenario in Table 3, ensuring consistency between the desired terminal speed and the proper-time dynamics of the wall.
Scenario $\eta\,[M_{p}^{4}]$ $\tau_{\rm term}\,[M_{p}^{-1}]$ $R_{\rm fin}\,[M_{p}^{-1}]$ $N_{\rm tot}^{\rm(int)}$
$\delta=10^{-8}$ $8.300019459820\mathrm{e}{-16}$ $7.069863\mathrm{e}{+07}$ $7.208700\mathrm{e}{+07}$ $1.052063\mathrm{e}{+08}$
$\delta=10^{-9}$ $624696593960\mathrm{e}{-16} 2.2$$235687\mathrm{e}{+08}$ $2.320814\mathrm{e}{+08}$ $3.530221\mathrm{e}{+09}$
$\delta=10^{-10}$ $8.300019398190\mathrm{e}{-17}$ $7.069863\mathrm{e}{+08}$ $7.147002\mathrm{e}{+08}$ $1.136138\mathrm{e}{+11}$
Table 3: For each ultra-relativistic deficit $\delta=1-v_{\rm term}$ we evolve the spherical wall with proper acceleration (7.4) and kinematics to the time $\tau_{\rm term}=\sigma/\eta$ where the terminal velocity is reached. The proper acceleration $\alpha(\tau)$ directly feeds the instantaneous occupancy (7.6); the total yield $N_{\rm tot}^{\rm(int)}$ follows from (7.8). We note that we took into account only the Higgs particles produced at rest, i.e. the $k=0$ mode. The number of particles produced in all momentum modes are certainly higher than shown here.
The results show that as the radius grows friction balances the drive, with the system approaching terminal velocity on the timescale $\tau_{\rm term}$; and the integrated yield is dominated by the combination of the exponentially sensitive $N_{k=0}(\alpha)$ and the rapidly growing area factor $4\pi R^{2}$, which becomes substantial by the time $\tau\sim\tau_{\rm term}$ in the most ultra-relativistic cases.
These Higgs bosons will decay almost exclusively into Standard Model final states, with the dominant channels producing high‐energy photons and neutrinos. In Figures 6 and 7 below, we present the resulting photon and neutrino spectra, respectively, showing the differential number of particles per event as a function of their energy in Planck‐mass units. These distributions encapsulate both the initial zero‐momentum mode occupation and the volume amplification that occurs during the ultra‐relativistic expansion of each bubble wall. In more intuitive terms, even though only a handful of quanta start life in the bubble wall’s zero mode, the rapid expansion amplifies them into a signal of Higgs decays—potentially producing observable flashes of photons and bursts of neutrinos in a cosmological first‐order transition.
Figure 6: The spectrum of photons generated by true‐vacuum bubbles via Higgs decay. The three curves correspond to Higgs particles produced with different terminal wall velocities, characterized by $\delta=10^{-8}$, $10^{-9}$, and $10^{-10}$, respectively, as summarized in Table 3.
Figure 7: The spectrum of neutrinos generated by true‐vacuum bubbles via Higgs decay. The three curves correspond to Higgs particles produced with different terminal wall velocities, characterized by $\delta=10^{-8}$, $10^{-9}$, and $10^{-10}$, respectively, as summarized in Table 3.
8 Photon or neutrino signal lead time
If expanding bubble walls propagate with a terminal velocity slightly below the speed of light, secondary particles such as photons or neutrinos emitted from Higgs decays will reach distant observers in advance of the wall itself. In this sense, they may act as an early “doomsday” warning signal. We now quantify the expected delay.
8.1 Cosmological inputs
[…]
$H_{0}$ $=\frac{67.4\times 10^{3}\;\mathrm{m\,s^{-1}}}{3.0857\times 10^{22}\;\mathrm{m}}\;\approx\;2.184\times 10^{-18}\;\mathrm{s^{-1}},$ (8.1)
$c$ $=2.99792458\times 10^{8}\;\mathrm{m\,s^{-1}},\qquad\Omega_{m}=0.315,\qquad\Omega_{\Lambda}=0.685.$ (8.2)
The Hubble rate evolves with redshift as
$H(z)=H_{0}\sqrt{\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}}.$ (8.3)
The corresponding comoving distance is
$\chi(z)=\int_{0}^{z}\frac{c}{H(z^{\prime})}\,dz^{\prime},$ (8.4)
while the photon travel time as measured by a comoving observer is
$t_{\gamma}(z)=\int_{0}^{z}\frac{dz^{\prime}}{(1+z^{\prime})\,H(z^{\prime})}.$ (8.5)
For illustration, we choose a source at proper distance
$D_{\rm ly}=1.0\times 10^{8}\;\mathrm{ly},\quad D_{m}=D_{\rm ly}\times 9.4607\times 10^{15}\;\mathrm{m},$
and determine the corresponding emission redshift by solving
$\chi\!\left(z_{\rm em}\right)=D_{m},$ (8.6)
which gives $z_{\rm em}\simeq 0.0069$.
8.2 Arrival delay for bubble walls
Consider now a bubble expanding at velocity
$v=(1-\delta)c,\qquad\delta\ll 1.$ (8.7)
In comoving coordinates the wall trajectory satisfies $d\chi/d\eta=1-\delta$, so the arrival delay relative to photons can be expressed in full generality as
$\Delta t(z)=\frac{\delta}{1-\delta}\int_{0}^{z}\frac{dz^{\prime}}{H(z^{\prime})}.$ (8.8)
This expression is valid at arbitrary redshift, and naturally reduces to the Minkowski estimate at low $z$. Expanding $H(z)\simeq H_{0}$ for $z\ll 1$ yields
$\Delta t\;\simeq\;\delta\,\frac{D}{c},$ (8.9)
with $D$ the proper distance today. This approximation suffices for the illustrative case $D=10^{8}$ ly ($z\simeq 0.007$).
Velocity Deficit $\delta$ Distance (ly) Time Delay
$\times 10^{-8} 1.0×10$$\times 10^{8}$ 1 yr, 0 d, 5 h, 58 m, 18.35 s
$\times 10^{-9} 1.0×10$$\times 10^{8}$ 0 yr, 36 d, 12 h, 35 m, 49.84 s
$\times 10^{-10} 1.0×10$$\times 10^{8}$ 0 yr, 3 d, 15 h, 39 m, 34.98 s
Table 4: Arrival delays for bubble walls with subluminal deficits $\delta$ relative to photons over a distance of $1.0\times 10^{8}$ light years, using the flat-space approximation $\Delta t\simeq\delta D/c$. For higher redshifts, Eq. (8.8) must be used.
Even a minute shortfall from luminal expansion produces a tangible lead time: photons or neutrinos emitted in Higgs decays would arrive days to months ahead of the wall itself. Such precursors could in principle serve as advance warnings of a cosmological true vacuum decay event. The three cases shown in Table 4 are purely illustrative; other choices of $\delta$ and redshift are straightforwardly evaluated using Eq. (8.8).
9 Conclusions
The work presented here was motivated by an interplay between the LHC collider data and recent theoretical progress in the field theory in curved spacetime. The standard model precision data coming from the LHC indicate that we currently live in a false vacuum. While the LHC data indicate that the lifetime of the false vacuum is safely large, it has been shown that small primordial black holes can serve as the phase transition catalysers and significantly increase the transition probability. In fact, it might be possible that the bubbles of the true vacuum already exist in our visible universe. One of the signatures of the existence of such bubbles would be copious production of the heavy Higgs particles due to the vacuum mismatch inside and outside of the bubble. We first calculated the spectrum of Higgs particles produced in this way using the standard field theory techniques. We then used the HDECAY event generator to simulate their decay. We finally used Pythia to hadronize the decay products and obtain the photon and neutrino energy spectra as a long-range signature of the bubble production.
The bubble wall’s velocity as seen by an external observer typically evolves in time as $v_{b}\sim t/\sqrt{t^{2}+R_{0}^{2}}$. Thus, for $t\gg R_{0}$, this velocity is almost the speed of light. However, it is never exactly equal to the speed of light. In fact, detailed calculations indicate that the bubble wall velocity might be limited to well below the speed of light [44]. In addition, the bubble wall can be viewed as a highly coherent state of a large number of the Higgs quanta. As such, it will strongly interact with virtually all the standard model particles. In the early universe, it will interact with the surrounding plasma, while at late times it will interact with other bubbles, stars, planets, interstellar and intergalactic gas, and finally even with the particles it creates itself. These interactions would further slow down the propagation of the wall. It is thus not very unreasonable to expect that these photons and neutrinos should be able to reach us before the bubble wall. No matter if this time delay is $10$ seconds or three years, if our gamma ray and neutrino observatories ever detect spectra corresponding to Figures 6 and 7, they might in fact represent the signals of the doomsday.
Since the particle production mechanism we relied on here resembles the Unruh effect, which in turn requires a non-zero proper acceleration of the wall, we included only particles produced until the terminal velocity is reached. After that, some other mechanisms might keep producing particles, for example because huge energy is dumped into the environment due to friction.
In broader context, a late time first‐order electroweak transition is expected to leave a multi‐messenger imprint— possibly a stochastic gravitational‐wave background, subtle distortions in the CMB and 21-cm line, and potential signals in gamma-ray, cosmic-ray and pulsar-timing data—that upcoming facilities are well poised to detect. These direct and indirect channels provide the most robust avenues for detecting the bubbles of the true vacuum. In a follow-up study, we will explore the detectability and present detailed forecasts for both current and next-generation observatories.
We note that the crucial assumption here is that there is no new physics beyond the standard model. New physics could potentially modify the Higgs potential and render our vacuum stable. Another possibility is that small primordial black holes do not exist. Since they are necessary catalysers, their absence in an appropriate mass range [18] would perhaps invalidate this doomsday scenario.
Acknowledgments
The authors are grateful to D. Wackeroth and C. Williams for their help with event generators and very useful comments on the paper. D.C. Dai is supported by the National Science and Technology Council (under grant no. 111-2112-M-259- 016-MY3). AS and DS are partially supported by the US National Science Foundation, under the Grant No. PHY-2310363.
10 Appendix-A
We derive here the magnitude of the proper acceleration of the bubble wall described by the equation $r^{2}-t^{2}=R_{0}^{2}$.
The bubble wall follows a hyperbolic trajectory in $(t,r)$ spacetime, which can be parametrized using proper time $\tau$ (the time experienced by an observer on the bubble wall). Note that this $\tau$ is not the same as the Euclidean time $\tilde{\tau}$ in Section 4. The proper parametrization is:
$t=R_{0}\sinh\left(\frac{\tau}{R_{0}}\right),\quad r=R_{0}\cosh\left(\frac{\tau}{R_{0}}\right).$
This satisfies the equation $r^{2}-t^{2}=R_{0}^{2}$, as:
$R_{0}^{2}\cosh^{2}\left(\frac{\tau}{R_{0}}\right)-R_{0}^{2}\sinh^{2}\left(\frac{\tau}{R_{0}}\right)=R_{0}^{2}.$
The 4-velocity $u^{\mu}$ of the bubble wall is the derivative of the coordinates with respect to proper time
$u^{\mu}=\left(\frac{dt}{d\tau},\frac{dr}{d\tau}\right)=\left(\cosh\left(\frac{\tau}{R_{0}}\right),\sinh\left(\frac{\tau}{R_{0}}\right)\right).$
Thus, the norm of the 4-velocity is
$u^{\mu}u_{\mu}=\left(\frac{dt}{d\tau}\right)^{2}-\left(\frac{dr}{d\tau}\right)^{2}=\cosh^{2}\left(\frac{\tau}{R_{0}}\right)-\sinh^{2}\left(\frac{\tau}{R_{0}}\right)=1,$
confirming that $\tau$ is the proper time.
The 4-acceleration $a^{\mu}$ is the derivative of the 4-velocity with respect to proper time
$a^{\mu}=\frac{du^{\mu}}{d\tau}=\left(\frac{1}{R_{0}}\sinh\left(\frac{\tau}{R_{0}}\right),\frac{1}{R_{0}}\cosh\left(\frac{\tau}{R_{0}}\right)\right).$
The norm of the 4-acceleration (which gives the proper acceleration) is
$a^{\mu}a_{\mu}=\left(\frac{1}{R_{0}}\sinh\left(\frac{\tau}{R_{0}}\right)\right)^{2}-\left(\frac{1}{R_{0}}\cosh\left(\frac{\tau}{R_{0}}\right)\right)^{2}=-\frac{1}{R_{0}^{2}}.$
The magnitude of the proper acceleration is the absolute value of the norm
$|a|=\sqrt{|a^{\mu}a_{\mu}|}=\frac{1}{R_{0}}.$
Thus, in the absence of friction, the magnitude of the proper acceleration of the bubble wall is constant and given by
$\dfrac{1}{R_{0}}.$
Since the relationship between proper and coordinate acceleration is $a=\gamma^{3}a_{\rm coor}$, as the wall’s speed approaches the speed of light ($v\rightarrow 1)$, the coordinate acceleration approaches zero, while the proper acceleration remains constant.
11 Appendix-B
Validation of the Thin-Wall Approximation
  • The true Higgs background is a smooth kink of width $\ell_{\rm wall}\sim 1/\mu$, interpolating between $\varphi_{\rm false}$ and $\varphi_{\rm true}$ over $\xi=(r-R(t))/\ell_{\rm wall}\in[-\tfrac{1}{2},\tfrac{1}{2}]$. In the thin-wall limit one sends $\ell_{\rm wall}\to 0$ at fixed surface tension $\sigma$, or equivalently requires
    $\varepsilon\;=\;\frac{\ell_{\rm wall}}{R(t)}\;\ll\;1.$
    For our fiducial nucleation parameters, $\ell_{\rm wall}\approx 1/\mu\approx 1.4\times 10^{3}\,M_{\rm Pl}^{-1}$ and $R_{0}=2\times 10^{4}\,M_{\rm Pl}^{-1}$, giving
    $\varepsilon=\frac{1/\mu}{R_{0}}=\frac{1.4\times 10^{3}}{2\times 10^{4}}\approx 0.07\;\ll\;1,$
    so curvature effects across the wall thickness are negligible.
  • In this regime the smooth mass profile $m^{2}(r,t)=\mu^{2}\,f(\xi)$ may be replaced by a sharp Heaviside step:
    $m^{2}(r,t)\approx\begin{cases}0\,,&r<R(t),\\ \mu^{2}\,,&r>R(t).\end{cases}$
    Finite-thickness corrections enter only at $\mathcal{O}(\varepsilon)$, modifying the Bogoliubov coefficients by $\mathcal{O}(\ell_{\rm wall}\,\mu)\sim\mathcal{O}(1)$ prefactors, but they cannot turn the exponentially suppressed factor $\exp(-4\mu/a)$ into an $\mathcal{O}(1)$ effect when $a<\mu$.
  • An independent WKB/adiabaticity check further confirms this. We define
    $\epsilon\;=\;\frac{1}{\omega^{3}}\Bigl{|}\tfrac{d}{dt}\,m^{2}\Bigr{|},\quad\omega\sim\mu.$
    By the chain rule,
    $\frac{d}{dt}m^{2}\sim m\,\dot{R}\,\frac{dm}{d\xi}\,\frac{1}{\ell_{\rm wall}}\sim\mu\times 1\times 1\times\mu=\mu^{2},$
    and thus
    $\epsilon\sim\frac{\mu^{2}}{\mu^{3}}=\frac{1}{\mu}\;\ll\;1.$
    Since $\dot{R}\lesssim 1$, $m\approx\mu$, $\tfrac{dm}{d\xi}\sim\mathcal{O}(1)$ (in units of wall widths), and $\ell_{\rm wall}\sim 1/\mu$, whenever $\mu\gg a$ we have $\epsilon\ll 1$, so mode mixing and particle production remain exponentially suppressed, $\sim e^{-\mu/a}$.
  • We conclude that using a thin wall approximation in our context is a controlled approximation at the few-percent level ($\varepsilon\approx 0.07$), and that finite-thickness effects cannot restore significant continuous production when $a<\mu$.